In a 1905 paper that later won him a Nobel Prize, Albert Einstein postulated that the energy of light was proportional to its frequency. The constant of proportionality turned out to be Planck's constant h: Elight=hf. Using your previous results, and the equation given in Part C, find an expression for h in terms of experimentally determinable quantities. Express your answer in terms of the slope m and e.
Answer #1

E = eV.    ............................(1)

E = hc/wavelength..............(2)

From 1 & 2 we write,.        V = (hc/e)(1/wavelength)

V = hf/e

Now comparing with, V = mf - b

                           Therefore, m = h/e

Hence, h = me.         where, m is slope = h/e